You are here: Foswiki>Commissioning Web>Com20080425 (25 Apr 2008, NormCushing)Edit Attach

Commissioning Notes for 20080425 UT

Summary:

Observer: Dave Thompson; TO: Randy Grashuis; Technical Support: John Hill; SW Support: Chris, Michele, Stephen, Jose, Torsten, Titus, Norm

Accomplishments...

  • Off-Axis and On-Axis guiding are successful at PA of 0, 90, 180
  • Guided for both Position Angle and Parallactic Mode.
  • Guided for 44 minutes before manually stopping to perform other activities.
  • During a 15 minute block (while it was examined) of the 44 minutes of guiding the star on IRTC only shifted .3 arcsecs.

At the beginning of the night...

TCS configuration: To test guide offsets and PCS responses, PCS was modified to not apply guide corrections (simulated guiding). Problem was found with guide offsets not being inverse translated into what is needed by PCS. Discussions with David Terrett underway.

Later in the night...

PCS was unpatched (so it would respond to guide corrections) and position angle and parallactic mode guiding was attempted and achieved.

PCS was modified to properly process guide information for Bent Gregorian instruments and the above list of activities was achieved.

There are a lot of details to complete but basic GCS based guiding is working.

Software changes requested for Friday night ...

  • GCS - provide way to manually take WFS images while guiding.
  • GCS - provide way to update guiding hot spot while guiding.
  • PCS - Support guiding for prime and BG focus instruments.
  • MCSPU/CSQ - Support on source flags for Az/El/Rotator for on-instrument and on-telescope rotators.
  • IRTC - Properly return information for get multi-parameter IIF call.

Running Notes/Log (UTtimes):

Running PCS with Ignore Guide Corrections in place

Open before sunset, everything seems to be fine

02:18 preset failed...restart CSQ (GCSL needed to be running, but the preset was in TRACK mode)

02:40 M5_0444 11:49:03.296 14:34:18.40 GS5
  • GS5 @ 94.70 E 294.05 S, guider stages should be: -49.32 424.07
  • Rejected as out of range by GCS
  • Manually moved stages and confirmed star was there.

02:55 Uploading a new TCS.

03:22 Starting new TCS build.

03:29 Preset fails (Invalid buffer: RPCERROR)

03:30 Attempting simulated guiding. Preset issued off-axis, PA = 0. Guider stage moved to proper position. Problem encountered with OSS. Trouble shooting in progress.

04:21 Use WT10_277 GS=0 to stay to the east...

04:25 Attempting simulated guiding but having trouble after acquiring saturated star. After 2 images it stopped guiding. Trouble shooting in progress.

04:45 Attempting simulated guiding. Took 1 frame and stopped taking pictures. Trouble shooting shows PCS was not replying to GCS message. Fix being implemented.

05:04 Installed new PCS.

05:06 Attempting simulated guiding. GCS is guiding. Move North and star moved up.

X,Y are GCS coordinates of on-axis star in mm. OffRA,OffDec are PCS telescope offsets in RA/DEC arcsec which are manually induced guiding disturbances. CentX,Cent are the computed pointing corrections in PCS from PointingControl:getGuide.

Data: PA=0 Rot=15.8 AZ=139.8 EL=72.5
X=0.4 Y=-1.2 OffRA=0 OffDec=2 CentX=1e-6 CentY=-7e-6
X=0.4 Y=+0.2 OffRA=0 OffDec=0 CentX=1e-6 CentY=+4e-6

East movement gave GCSL died unexpectedly.

05:15 Attempting simulated guiding. Issue preset. Offset 2,0 (arcsec). Image of star moved to the left.

X=0.5 Y=-0.3 OffRA=0 OffDec=0 CentX=2e-7 CentY=3e-6
X=-0.8 Y=-0.3 OffRA=2 OffDec=0 CentX=-8e-6 CentY=2e-6

05:18 Attempting simulated guiding. Preset with PA=90. Offset 0,2 (arcsec). Image of star moved to the right. Move back to 0,0. Image moved to right(?). Move to 2,0. Star moved up.

Data: PA=90 Rot=97.7 AZ=149.5 EL=74.2
X=-0.5 Y=-0.5 OffRA=0 OffDec=0 CentX=2e-6 CentY=-3e-6
X=0.9 Y=-0.6 OffRA=0 OffDec=2 CentX=-7e-6 CentY=-3e-6
X=-0.6 Y=-0.6 OffRA=0 OffDec=0
X=-0.6 Y=-1.7 OffRA=2 OffDec=0 CentX=-6e-6 CentY=-1.2e-5

05:28 Attempting simulated guiding. Preset with PA=0. Slew to new target. Guider did not find the image. Exposure was low, .001 ms as documented in FITS header. Star 150 counts above background. GCS needs at least 200 counts above background.

05:30 Guide camera exposure changed to .1 ms

05:30 Added the Persson faint IR Standards to the default catalog. Printout included in the notebook. These are fainter (V~10-14 mag), so the AGW can integrate longer to average over the seeing.

Checking max exposure time. About half-well in 2s, but Persson Std saturated

Taking data for H-band flat:
  • 060617 + 00003_cube
  • 060719 + 00004_cube (offset 9.0 0.0 from 1st)
  • 060801 + 00005_cube (offset -9.0 0.0 from 1st)
  • 060854 + 00006_cube (offset 0.0 9.0 from 1st)
  • 060936 + 00007_cube (offset 9.0 0.0 from 1st) Need 2s darks!

Taking data for J-band flat:
  • 061445 + 00008_cube
  • 061603 + 00009_cube (offset 9.0 0.0 from 1st)
  • 061708+ 00010_cube (offset -9.0 0.0 from 1st)
  • 061752 + 00011_cube (offset 0.0 9.0 from 1st)
  • 061835 + 00012_cube (offset 0.0 -9.0 from 1st) Need 2s darks!

Checking open-loop tracking with 2s exposures (100x sequence)
  • 062705 + 00013_cube

05:37 Attempting simulated guiding. Preset with PA=0. Guider did not find the image. Exposure is .1 in FITS header. Star still at 150 counts above background. Cannot explain why this happened. Dave T determined the telescope was out of focus. Will try again after telescope is in focus.

05:54 Attempting simulated guiding. Preset with PA=0. Guider found the star and started guiding. Offset of 0,2. Star moved up. Offset 2,0. Star moved left.

Data: PA=0 Rot=120.7 AZ=28.0 EL=66.9
X=0.06 Y=-0.04 OffRA=0 OffDec=0 CentX=-1e-6 CentY=-6e-7
X=-0.2 Y=-1.3 OffRA=0 OffDec=2 CentX=-3e-6 CentY=-1.1e-5
X=-1.5 Y=-0.5 OffRA=2 OffDec=0 CentX=-1.2e-5 CentY=-3e-6

05:56 Attempting simulated guiding. Preset with PA=90. Guider found the star. Offset 0,2. Star moved right. Offset 2,0. Star moved up. Image is 1.25 arcsec FWHM.

Data: PA=90 Rot=-147.3 AZ=26.8 EL=67.2
X=-0.03 Y=0.01 OffRA=0 OffDec=0 CentX=-6e-7 CentY=-1e-7
X=1.2 Y=-0.2 OffRA=0 OffDec=2 CentX=9e-6 CentY=0
X=-0.07 Y=-1.15 OffRa=2 OffDec=0 CentX=-2e-6 CentY=-9e-6

06:00 Spending time examining PCS log files to understand how PCS is responding to guide corrections. (see data on CentX, CentY in radians reported above)

Conclusion is that pointing kernel is scaling the X,Y corrections in mm directly to AZ/EL corrections in radians without bothering to correct them for the rotator angle. Now we have to understand why that is happening........

Running PCS in Parllactic Mode and honoring guide corrections

06:35 On to some GCS tests...

Tonight we were able to guide on an off-axis source in PA mode for ~50 minutes!

2s Darks for the sky flat data (30x sequences):
  • 120332 + 00047_cube
  • 120449 + 00048_cube

For the record - this camera has huge dark current. The "sky" frames I took tonight are 8000DN of dark and ~80DN of sky photons. I'll try again tomorrow evening before the sun sets.

07:52 PCS is being unpatched to allow guiding so we can test parallactic angle mode guiding on-axis.

08:04 Parallactic Angle Preset. Star appears in IRTC image. Rotator Angle is 340, RA=4.271, Dec=?

08:08 Issue Guiding in Parallactic Mode Preset. Guide probe moved to 600.97 and GCS reported probe position was out of range so guiding failed. GCS should be investigated but this is also being investigated as a PCS preset initialization issue.

08:14 Issue Guiding in Parallactic Mode, PA=0. Moving to fainter star. Setting guide exposure to 1s. Guide image was driven out of the field of view.

08:18 Issue Guiding in Parallactic Mode, PA=180 and trying again. Guiding was achieved. Rotator Angle 160:52.

08:23 Attempt stopping guiding and than closing the loop. It failed saying no guide star in the list.

08:30:30 GCS died unexpectantly. GCS restarted at 1:41

08:43 Will issue new preset for Guiding in Parallactic Mode, PA=180, tilted mirror to put WFs camera in the right place. GCS did not register its rpc. Attempted to netconfig stop GCSL it said, GCS constructor finished at 1:43:56 (local time), registering commands, registration complete, starting GCS thread. GCS is hung.

08:47 Restarted GCSL

08:46 Will issue new preset for Guiding in Parallactic Mode, PA=180, tilted mirror to put WFs camera in the right place. GCSL died 6 seconds after doing an _status at 01:48:12 (local time).

08:50 Restarted GCSL, issued Preset, guiding!!!

Running PCS with proper guide handling for BG instruments.

09:15 PCS upgraded with guiding fix.

09:21 Attempt on-axis guiding, PA=90. Moving pointing offset to move guide star was corrected by GCS and PCS to keep star centered and guiding going!!!

09:28 Attempt off-axis guiding. PA=14.34, Position Angle Mode, 235 arcsecs off axis, rmag of 10, exposure set to 250 ms. It works!!! Centroid is slightly left of center in all images. Pointing offsets to throw it off resulted in guiding bringing the star back to desired position.

09:38 Attempt off-axis guiding. PA=242, Position Angle Mode.

Updated exposure time during guiding from cmd line commands to 1000 ms. It took 180 seconds to take affect. Guiding stopped. Not sure if related.

09:47 Attempt off-axis guiding. PA=238.03, Position Angle Mode.

09:48 Noticed that MCSPU log file mcs.log has not updated in 4 hours. MCSPU was out of disk space. Impressive that it was able to run for 4 hours without disk space. We deleted some 2-day old core files to make a small amount of space. NEEDS ATTENTION IN THE MORNING!

09:56 GCS has been hung for many minutes in a preset command. Trying to stop and restart GCSL, but even several "netconfig kill GCSL" commands on lbtmu107 have trouble killing it.

10:05 Preset again, and guiding on the same star.

10:49 Still guiding all this time. Stopped on purpose.

11:00 - 11:30 (04:00 - 04:30 local time) Polynomials keep flowing!!! Cron jobs had been set to run at 4:02 for dailies and 4:22 for weeklies and 4:42 for monthlies. Those have been changed to now run during the 10 AM hour (local time).

Polynomials did drop out several times earlier in the evening. Suspect the cause to be a lot of PCS I/O for debugging and other developer activities on the same server as that of PCS. Speculation is removing I/O will make an improvement. Noticed the machine is running with ~45% waiting.

12:19 IRTC TE cooler shut off and logged out of the GUI. AGW homed but left running.

-- NormCushing and DavidThompson - 25 Apr 2008
Topic revision: r6 - 25 Apr 2008, NormCushing
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback